OpenStudy (anonymous):

20. You join a paperback book club for $6 and then spend $3 per book. Y = 3x + 6. Explain how to graph the equation, using the y-intercept and slope. Be sure to tell where you place the first point and how to find the location of the second point

OpenStudy (goten77):

hmm well the y intercept is when x=0 so y= 3(0) +6 y=6 is where the graph will cross the y axis that coordinate is (0,6) the slope is 3 meaning move up 3 then across 1 also the x intercept is when y=0 so then we have 0=3x+6 -6=3x -2=x so the x intercept is at the coordinate (-2,0) and now we know how the function behaves and where it crosses both axes's